# MARS: A Multi-Agent Framework Incorporating Socratic Guidance for Automated Prompt Optimization

### v1

@@ Abstract
The basic question-answering format of large language models involves inputting a prompt and receiving a response, and the quality of the prompt directly impacts the effectiveness of the response. Automated Prompt Optimization (APO) aims to break free from the cognitive biases of manually designed prompts and explores a broader design space for prompts. However, existing APO methods suffer from limited flexibility of fixed templates and inefficient search in prompt spaces as key issues. To this end, we propose a Multi-Agent framework IncorpoRating Socratic guidance (MARS) [^1] The code and experiment datasets are available at https://github.com/exoskeletonzj/MARS, which utilizes multi-agent fusion technology for automatic planning, with gradual continuous optimization and evaluation. Specifically, MARS comprises seven agents, each with distinct functionalities, which autonomously use the Planner to devise an optimization path that ensures flexibility. Additionally, it employs a Teacher-Critic-Student Socratic dialogue pattern to iteratively optimize the prompts while conducting effective search. We conduct extensive experiments on various datasets to validate the effectiveness of our method, and perform additional analytical experiments to assess the model’s advancement as well as the interpretability.
MARS: A Multi-Agent Framework Incorporating Socratic Guidance
for Automated Prompt Optimization
Jian Zhang1,3, Zhangqi Wang1, Haiping Zhu1, Jun Liu1, Qika Lin2thanks: Corresponding author, Erik Cambria3∗ 1Xi’an Jiaotong University 2National University of Singapore 3Nanyang Technological University zhangjian062422@stu.xjtu.edu.cn, qikalin@foxmail.com, cambria@ntu.edu.sg
1 Introduction
Large language models (LLMs) such as GPT-4 (Achiam et al., 2023) and Deepseek-R1 (Guo et al., 2025) provide robust support for thousands of natural language processing tasks. By providing a natural language prompt that includes instructions and a task description, LLMs can quickly adapt and respond Lin et al. (2025). Consequently, the quality of the prompt is of critical importance, leading to wide interest in Automated Prompt Optimization (APO) (Pryzant et al., 2023). As shown in Figure 1, we provide LLMs with three different inputs for the word sorting task: a zero-shot prompt, a Chain of Thought (CoT) prompt, and our optimized prompt. The responses are produced in a markedly distinct way. Specifically, the zero-shot prompt incorrectly identifies the alterate as the more common word alternate. However, the task requires faithfully preserving the given sequence of words rather than correcting them. With the CoT prompt, the sorting remains incorrect because the LLM does not fully grasp the sorting task and the word sequence. In contrast, our optimized prompt produces the correct answer. This is because our prompt includes specific requirements, such as maintaining the original letter casing and specifying the sorting method.
@@ 1 Introduction
Thus, it is evident that APO produces better results when accomplishing the task. Some recent studies (Zhou et al., 2022; Xu et al., 2023; Wang et al., 2023) optimize prompts by generating multiple candidates and employing various search strategies, while others (Yang et al., 2024a; Ye et al., 2023) focus on crafting sophisticated meta prompts for optimization. However, two major issues persist in this task: limited flexibility of fixed templates and inefficient search in prompt spaces.
The first issue is the limited flexibility of fixed templates. In previous research (Yang et al., 2024a; Ye et al., 2023), meta prompts are fixed optimization templates with limited flexibility, typically relying on predefined templates. Unlike tasks such as event extraction (Zhang et al., 2024b) and text2symbol (Xu et al., 2024, 2025), where fixed templates can be used to complete tasks, these templates could not be dynamically adjusted to meet the varying needs of different tasks, thus limiting the effectiveness of the method. This leads to poor performance when handling diverse tasks or complex scenarios, as the fixed templates may introduce biases or fail to optimize effectively, especially when tasked with addressing the varied requirements of different tasks.
The second issue is inefficient search in prompt spaces. Some methods (Zhou et al., 2022; Xu et al., 2023; Wang et al., 2023) using the generation-search strategy generate a cluster of prompts within the prompt space and then optimize them using search strategies within that cluster. However, this is a local optimization approach, focusing only on optimizing within the pre-generated set of prompts. As a result, the search in the entire prompt space may be incomplete, leading to ineffective search results and limited overall optimization of prompts.
To this end, we propose a Multi-Agent framework integRating Socratic guidance (MARS). Specifically, we construct a multi-agent architecture that includes a Planner agent and six task-specific agents for APO. First, to tackle the first issue, we employ the Planner agent to autonomously plan optimization paths for different tasks, ensuring that each task follows its own optimization path. Next, due to the advantages of Socratic questioning Elder and Paul (1998) in promoting critical thinking and stimulating autonomous learning, we employ a Teacher-Critic-Student Socratic guidance dialogue pattern to iteratively optimize the prompts while ensuring interpretability. This module enables MARS to search across the entire prompt space and, through feedback, continuously narrow down the prompt space, effectively searching for the optimal prompt. Finally, we validate the optimized prompts in the Target agent, iterating until the optimal prompt is found. We conduct extensive experiments across multiple general tasks and domain-specific datasets to validate the effectiveness of our designed model and explore the interpretability of the optimization process and results. In summary, the contributions of this paper are in following three folds:
$\bullet$ We propose a multi-agent architecture for APO that autonomously plans optimization steps, effectively addressing the issue of limited flexibility of fixed templates. To the best of our knowledge, this is the first attempt in the field.
$\bullet$ We introduce a Teacher-Critic-Student Socratic guidance dialogue pattern for the APO task, wherein the Teacher employs Socratic questioning, the Critic evaluates the quality of those questions, and the Student undergoes optimization to achieve the optimal prompt.
@@
$\bullet$ Extensive experiments on both general tasks and domain-specific task datasets demonstrate the effectiveness and superiority of the MARS framework. Additional analytical experiments further validate the interpretability of the optimization process and results.
2 Methodology
In this section, we provide a detailed introduction to multi-agent framework incorporating Socratic guidance. We base our design on LLMs as intelligent agents capable of performing various functions. The MARS framework autonomously plans the optimization path and employs a Socratic guidance dialogue pattern to flexibly plan the optimization path and address the issue of inefficient search in the prompt space. The model architecture is shown in Figure 2. In Section 2.1, we introduce the task definition. In Section 2.2, we present the multi-agent framework for APO. In Section 2.3, we discuss the Teacher-Critic-Student Socratic guidance dialogue pattern. And in Section 2.4, we cover effect validation and iterative optimization.
Figure 2: The overall architecture of the MARS model. It consists of seven LLM agents. The Manager oversees the entire process, responsible for communication between agents and the allocation of speakers. The UserProxy receives input and the original prompt. The Planner formulates the APO plan based on the received input and task description. In the Socratic guidance dialogue pattern, the Teacher-Critic-Student system iteratively refines the prompts according to the Planner’s plan, with the final evaluation and recording done by the Target.
2.1 Task Definition
Following the definition in prompt optimization Zhou et al. (2022), when using the given target agent $\mathcal{M}_{tar}$ as the task model, the goal of the APO task is to start from the original text prompt $p_{0}$, progressively optimize it, and ultimately obtain the text prompt $p^{*}$ that achieves the best performance on the given dataset $D=\{(x,y)\}$. We will select a few examples from ${D_{train}}$ to optimize the prompts and finally use ${D_{test}}$ for evaluation. The formal definition of this process is as follows:
[…]
where $\mathcal{M}_{tar}(x;p)$ is the response obtained when the input $x$ and the prompt $p$ are fed together into the target agent, and $f$ is the function to measure the model’s performance (e.g., accuracy).
2.2 Multi-Agent Framework
Figure 2 illustrates our multi-agent MARS architecture incorporated with Socratic guidance. This architecture includes essential modules as well as several innovative modules for multi-agent systems. This section first introduces three agents: Manager, UserProxy, and Planner, with the other agents to be introduced in Sections 2.3 and 2.4.
Manager. The group chat Manager acts as the administrator, overseeing the entire process and allocating speaking rights to ensure effective coordination among the agents.
UserProxy. The UserProxy agent serves as the input receiver, responsible for accepting and processing external inputs, and providing informational support to other agents, thus facilitating the smooth operation of the entire system.
Planner. The Planner agent is responsible for planning the task based on the input, breaking it down into several optimization sub-steps. In our architecture, the entire planning process is handled by the Planner agent, which fully leverages its extensive knowledge and reasoning capabilities. This allows the Planner agent to provide clear guidance for the subsequent optimization process, ensuring that each sub-step effectively advances the task. The optimization will then proceed step-by-step according to the path outlined by the Planner, optimizing the final output while ensuring task completion. The planned steps $\mathbf{st}$ is shown as following:
$\mathbf{ST}[st_{1},st_{2},\dots,st_{n}]=\underset{x\in D_{\text{train}}}{\mathcal{M}_{\text{plan}}}(x;p_{0}),$ (2)
where $st_{i}$ is the plan for the current $i$-th step, $\mathcal{M}_{plan}$ is Planner agent, and $\mathbf{ST}[...]$ is the optimization steps planned for the current task.
2.3 Socratic guidance Dialogue
Previous APO methods (Zhou et al., 2022; Xu et al., 2023; Yang et al., 2024a; Ye et al., 2023) passively rely on LLMs to optimize prompts, either by generating multiple candidates and searching or by refining prompts using meta prompts. However, these approaches do not provide insight into why a particular optimization is effective or why the optimized prompt performs better.
Constructivist learning theories (Hein, 1991) emphasize two fundamental principles of Socratic teaching (Elder and Paul, 1998; Park et al., 2024; Liu et al., 2024a). First, it is inherently dialogic, with the learning process facilitated through interactions and conversations between teachers and students. Second, it employs probing questions to actively engage students, fostering independent thinking and encouraging them to find answers on their own. Inspired by Socratic teaching, we introduce an active optimization approach: the Teacher agent continuously proposes open-ended questions based on the planned optimization steps, the Critic agent evaluates the quality of these questions, and the Student agent engages in the dialogue, actively participating in the APO process. This approach encourages the Student to express ideas, challenge assumptions, think independently, and ultimately produce the optimal prompt. Algorithm 1 presents the detailed design of the Teacher-Critic-Student Socratic guidance dialogue algorithm, clearly illustrating its core logic and implementation steps.
@@ 2.3 Socratic guidance Dialogue
As shown in Figure 3, the Teacher-Critic-Student Socratic guidance dialogue pattern is used to iteratively optimize the prompt, involving three key agents:
Teacher. Based on the sub-steps planned by the Planner, the Teacher agent uses a Socratic style to guide the Student in solving the problem of each sub-step. There are three main requirements for the Teacher agent: First, it should not directly provide the answer, but instead guide the Student to think through the solution process using probing questions. Second, it should explain the content of each sub-step and the knowledge involved. Third, the Teacher agent must maintain a consistent teaching style throughout to ensure continuity in the instruction. The question $q[st_{i}]$ for the $i$-th step is as follows:
$q[st_{i}]=\underset{st_{i}\in\mathbf{ST}}{\mathcal{M}_{\text{teacher}}}(st_{i};p_{i-1}),$ (3)
where $p_{i-1}$ is the prompt optimized in the ${(i-1)}$-th round, and ${\mathcal{M}_{\text{teacher}}}$ is the Teacher agent.
Algorithm 1 MARS Procedure
1:  Input: Task Dataset $\mathcal{D}$, Original Prompt $p_{0}$
@@
3:  Manager: Initialize the APO task
4:  UserProxy: Receive the input $x\in D_{train}$, prompt $p_{0}$
5:  Planner: Plan optimization Steps $\mathbf{ST}={\{st_{i}\}}$
6:  for iteration $=1$ to $I$ do
7:     for each step $st_{1},st_{2},...,st_{n}\in\mathbf{ST}$ do
8:        Teacher: generate socratic question $q[st_{i}]$ of $st_{i}$ review the student’s answer $p_{i-1}$
9:        Critic: assess the Socratic style of Teacher’s Question $q[st_{i}]$
10:        if TRUE then Continue
11:        else Teacher: adjust the question $q[st_{i}]$
12:        end if
13:        Student: optimize the prompt $p_{i}$
@@
15:     Target: Evaluate $(x,y)\in D_{test}$ with $\mathcal{M}_{target}(x;p_{n})$
16:  end for
17:  return  Best prompt $p^{*}$
Critic. As a key component, the Teacher agent plays a crucial role in determining the effectiveness of the student’s learning through the quality of its questions. Therefore, we introduce the Critic agent as a supervisor. If the Critic agent judges that the Teacher agent’s questions do not adhere to the Socratic style, it provides feedback for revision. This judge-revise cycle continues until the questions meet the required standard, at which point they are passed on to the Student agent.
Models B.E D.QA F.F. G.S. R.N. S.U. C.B. C.M. E.E. W.H. H.A. M.T. Avg.
@@
Student. The Student agent generates responses based on the questions posed by the Teacher, gradually working through each step of the optimization process. By actively engaging in the dialogue, the Student agent deepens its understanding of the key elements of each sub-step. Throughout this process, the Student agent adjusts its thought process in response to the Teacher’s guidance and the Critic’s feedback, ultimately arriving at the final prompt for this optimization cycle. The Student’s independent thinking and active participation are crucial in ensuring that the resulting prompt is both logically sound and effective.The prompt $p_{i}$ for the $i$th step is as follows:
$p_{i}={\mathcal{M}_{\text{student}}}(q[st_{i}];p_{i-1}).$ (4)
Specific examples and analysis of the Socratic guidance dialogue are provided in Appendix F.
2.4 Effect Validation
Target. The prompt optimized through the previous steps will be evaluated on the test dataset $D_{test}$ of the Target agent. Based on the results of the first round of optimization, a second round of optimization will be carried out, followed by another evaluation on $D_{test}$. This iterative process continues until we reach the number of iterations we have set (e.g., 10 times). The evaluation $p_{i}$ for the $i$-th step is as follows:
$y_{pred}=\underset{{(x,y)\in{D_{test}}}}{\mathcal{M}_{\text{tar}}}(x;p_{i}).$ (5)
Where $y_{pred}$ is the predicted result of $p_{i}$ under the target agent. The accuracy is calculated by comparing $y_{pred}$ with the gold label $y$.
3 Experiments
In this section, we conduct extensive experiments on 12 general task datasets and 5 domain-specific datasets. The baseline methods and evaluation metrics are introduced in Section 3.1, while the main experimental results are presented in Sections 3.2. A detailed analysis of efficiency is provided in Section 3.3.
3.1 Datasets and Baselines
Tasks and Datasets
We selected a total of 17 datasets from both general and domain-specific tasks for extensive experimentation, which include: 6 BBH (Suzgun et al., 2022) general tasks, 6 MMLU (Wang et al., 2024b) general tasks, 3 C-Eval (Huang et al., 2024) Chinese domain tasks, 1 LSAT-AR (Zhong et al., 2023) law domain task, and 1 GSM8K (Zhang et al., 2024a) math domain task. More detailed descriptions of the tasks and datasets can be found in the Appendix A. The abbreviations of the tasks in Table 1 and Table 2, along with their full names and dataset descriptions, are provided in Table 4. The experiment settings and baselines are introduced in Appendix B.
3.2 Main Results

### v2

# MARS: Multi-Agent Adaptive Reasoning with Socratic Guidance for Automated Prompt Optimization

@@ Abstract
Large language models (LLMs) typically operate in a question-answering paradigm, where the quality of the input prompt critically affects the response. Automated Prompt Optimization (APO) aims to overcome the cognitive biases of manually crafted prompts and explore a broader prompt design space. However, existing APO methods often suffer from rigid template structures and inefficient exploration in the prompt space. To this end, we propose a Multi-Agent Adaptive Reasoning with Socratic guidance framework (MARS) for APO. MARS consists of five complementary agents and formulates the optimization process as a Partially Observable Markov Decision Process (POMDP), enabling adaptive prompt refinement through explicit state modeling and interactive feedback. Specifically, a Planner agent generates flexible optimization trajectories, a Teacher-Critic-Student triad engages in Socratic-style dialogue to iteratively optimize the prompt based on pseudo-gradient signals in the text space, and a Target agent executes the prompt in downstream tasks to provide performance feedback. MARS integrates reasoning, feedback, and state transition into a unified hidden-state evolution process, improving both the effectiveness and interpretability of optimization. Extensive experiments on multiple datasets demonstrate that MARS outperforms existing APO methods in terms of optimization performance, search efficiency, and interpretability.
Codehttps://github.com/exoskeletonzj/MARS
Introduction
Large language models (LLMs) such as GPT-4 (achiam2023gpt) and Deepseek-R1 (guo2025deepseek) provide robust support for thousands of natural language processing tasks (ijcai2025p772). By providing a natural language prompt that includes instructions and a task description, LLMs can quickly adapt and respond (lin2025self) (shentie2025) (yan2025mur). Consequently, the quality of the prompt is of critical importance, leading to wide interest in Automated Prompt Optimization (APO) (pryzant2023automatic). As shown in Figure 1, we provide LLMs with three different inputs for the word sorting task: a zero-shot prompt, a Chain of Thought (CoT) prompt, and our optimized prompt. The responses are produced in a markedly distinct way. Specifically, the zero-shot prompt incorrectly identifies the alterate as the more common word alternate. However, the task requires faithfully preserving the given sequence of words rather than correcting them. With the CoT prompt, the sorting remains incorrect because the LLM does not fully grasp the sorting task and the word sequence. In contrast, our optimized prompt produces the correct answer. This is because our prompt includes specific requirements, such as maintaining the original letter casing and specifying the sorting method.
@@ Introduction
Figure 2: Comparison of APO strategies. Generation-search and meta-prompt. Multi-Agent Adaptive Reasoning enables dynamic, collaborative reasoning. Right: With GPT-4o, MARS outperforms all baselines on three benchmarks.
Thus, it is evident that APO can lead to improved performance in downstream tasks. As shown in Figure 2, recent studies (zhoularge; xureprompting; wang2023promptagent) have explored prompt optimization by generating multiple candidates combined with diverse search strategies, while others (Yang0LLLZC24; ye2023prompt) focus on designing sophisticated meta-prompts to guide optimization. Despite these advances, two key issues remain: the limited flexibility of fixed prompt templates, and the inefficiency of prompt space exploration.
The first issue is the limited flexibility of fixed templates. Prior works (Yang0LLLZC24; ye2023prompt) often rely on meta-prompts, which are predefined optimization templates that cannot be dynamically adapted to different tasks. Unlike domains such as event extraction (zhang2024semantic) or text-to-symbol generation (xu2023symbol), where fixed templates suffice due to the task’s structural regularity, APO requires more adaptability. Rigid templates may introduce biases or fail to capture task-specific nuances, resulting in suboptimal performance when applied to diverse or complex scenarios.
The second issue is the inefficiency of prompt space exploration. Several approaches (zhoularge; xureprompting; wang2023promptagent) adopt a generation-search strategy, where a set of candidate prompts is first generated and then refined using local search techniques. However, this approach typically performs only local exploration around the initial candidates, without sufficiently covering the broader prompt space. As a result, the optimization may converge prematurely or overlook better-performing prompts, limiting overall effectiveness.
To this end, we propose a Multi-Agent Adaptive Reasoning with Socratic guidance framework (MARS) for APO. MARS consists of five complementary agents and formulates the optimization process as a Partially Observable Markov Decision Process (POMDP), enabling adaptive prompt refinement through explicit state modeling and interactive feedback. Functionally, to address the first challenge, MARS introduces a Planner agent that generates task-specific optimization trajectories, allowing prompts to be flexibly adapted to diverse task requirements. To tackle the second challenge, MARS employs a Socratic-style Teacher-Critic-Student dialogue mechanism, which iteratively guides prompt refinement. This module enables effective exploration of the prompt space by simulating a gradient-inspired optimization process, while also promoting interpretability. The overall process is modeled as a POMDP, where the hidden state represents the latent reasoning state of the Student agent. Through multi-agent interactions and performance feedback from a Target agent, MARS approximates a pseudo-gradient trajectory in the discrete prompt space, progressively refining the prompt toward an optimal solution. Our contributions are three-fold:
$\bullet$ This work is the first to introduce a multi-agent architecture with POMDP modeling for APO. It proposes MARS, which enables hidden-state reasoning and adaptive planning through agent collaboration.
$\bullet$ A Teacher-Critic-Student Socratic dialogue mechanism is designed to enable interpretable, iterative prompt refinement via a gradient-inspired optimization trajectory.
@@
$\bullet$ We demonstrate the effectiveness and generalizability of MARS through extensive experiments on both general and domain-specific benchmarks, and validate the interpretability of its optimization process.
Figure 3: The overall architecture of the MARS model. It consists of five LLM agents. The Planner agent that autonomously generates task-specific optimization trajectories, and a Teacher-Critic-Student Socratic dialogue mechanism that iteratively refines prompts, with the evaluation and iterative refinement process guided by feedback from the Target agent.
Methodology
MARS comprises two main modules: (i) a high-level Planner that generates task-specific optimization trajectories, and (ii) a Teacher-Critic-Student triad that performs Socratic-style iterative refinement. The overall architecture is shown in Figure 3, with the complete workflow detailed in Algorithm 1. This section introduces: (1) the APO task and its POMDP formulation, (2) the Planner design, (3) the gradient-inspired Socratic dialogue mechanism, and (4) the evaluation-feedback loop via the Target agent.
Task Formulation and POMDP Modeling
Given a task-specific Target model $\mathcal{M}_{\text{tar}}$, the goal of APO is to iteratively refine an initial prompt $p_{0}$ to an optimal version $p^{*}$ that maximizes performance on a downstream dataset $D=\{(x,y)\}$. A training subset $D_{\text{train}}\subset D$ guides the optimization, while $D_{\text{test}}$ is used for evaluation. The objective can be formalized as:
$p^{*}=\arg\max_{p}\sum_{(x,y)\in D_{\text{test}}}f\left(\pi_{\text{tar}}(x;p),y\right),$ (1)
where $\pi_{\text{tar}}(x;p)$ denotes the model output conditioned on $x$ and prompt $p$, and $f$ is a task-specific metric (e.g., accuracy, BLEU).
To capture the sequential, partially observable nature of the optimization, we model APO as a Partially Observable Markov Decision Process (POMDP) defined by:
$\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\mathcal{O}\rangle,$
where: - $\mathcal{S}$: latent state space, representing the internal reasoning state of the Student agent; - $\mathcal{A}$: action space, comprising instructional signals (e.g., questions, critiques) from Teacher and Critic; - $\mathcal{T}\!:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$: transition dynamics, updating student states; - $\mathcal{O}\!:\mathcal{S}\rightarrow\mathcal{P}$: observation function, mapping hidden states to prompts; - $\mathcal{R}(s,a)=f(\pi_{\text{tar}}(x;\mathcal{O}(s)),y)$: reward function, based on performance of the generated prompt.
This formulation allows MARS to perform gradient-inspired prompt refinement in a partially observable, discrete text space. Via iterative multi-agent reasoning and feedback, the system progressively transitions from $p_{0}$ to $p^{*}$.
Optimization Trajectory Planning
As illustrated in Figure 3, MARS begins with a Planner agent that initiates the prompt optimization process.
Planner. Given task goal $g$, input $x\in D_{\text{train}}$, and initial prompt $p_{0}$, the Planner decomposes the optimization into a sequence of sub-goals:
$\mathbf{ST}=[st_{1},st_{2},\dots,st_{n}]=\pi_{\text{plan}}(g,x,p_{0}),$ (2)
where $\pi_{\text{plan}}$ is the planning policy that adaptively generates an optimization trajectory.
To formalize $\pi_{\text{plan}}$, we introduce a latent planning variable $z\in\mathcal{Z}$, and model trajectory generation as:
$\pi_{\text{plan}}(g,x,p_{0})=\arg\max_{\mathbf{ST}}\,\mathbb{E}_{z\sim q(z|g,x)}\left[\log P(\mathbf{ST}\mid z,p_{0})\right],$ (3)
where $q(z|g,x)$ captures task semantics, and $P(\mathbf{ST}\mid z,p_{0})$ models the trajectory conditioned on latent intent and initial prompt. This hierarchical formulation induces structured plans over latent space $\mathcal{S}$, guiding local agent decisions under global coherence and improving adaptability over static templates.
Socratic Prompt Refinement as Joint Policy Optimization
Prompt refinement in discrete language space presents unique challenges due to its non-differentiability, high variance, and semantic ambiguity. To address these issues, MARS employs a structured Socratic dialogue mechanism involving three collaborative agents—Teacher ($\pi_{\text{t}}$), Critic ($\pi_{\text{c}}$), and Student ($\pi_{\text{s}}$)—each fulfilling a complementary role in exploring and improving prompts through guided interaction. This framework transforms prompt optimization into an interpretable, policy-driven reasoning process grounded in pedagogical principles.
At each refinement step $i$, given a sub-goal $st_{i}\in\mathbf{ST}$, the Teacher proposes a Socratic-style question $q_{i}$ to stimulate reasoning, based on the prior prompt $p_{i-1}$. The Critic then assesses its clarity, relevance, and coherence, producing feedback $c_{i}$ to revise or validate the proposed direction. Finally, the Student responds by updating its internal state and generating a new prompt $p_{i}$. This process is formalized as:
$q_{i}$ $=\pi_{\text{t}}(st_{i},p_{i-1}),$ (4)
$c_{i}$ $=\pi_{\text{c}}(q_{i}),$
$p_{i}$ $=\pi_{\text{s}}((q_{i},c_{i}),p_{i-1}),$
$s_{i}$ $\sim\mathcal{T}(s_{i-1},(q_{i},c_{i})),\quad o_{i}=p_{i}.$
Each agent performs a partial update to the joint optimization process: Teacher drives semantic direction, Critic provides quality control, and Student synthesizes the final output.
Context-Aware Interaction.
To improve reasoning consistency and avoid step-wise myopia, each agent conditions not only on the current sub-goal and prompt, but also on the dialogue history $\mathcal{H}_{<i}=\{(q_{j},c_{j},p_{j})\}_{j=1}^{i-1}$. The full context-aware behavior is given by:
$q_{i}$ $=\pi_{\text{t}}(st_{i},p_{i-1},\mathcal{H}_{<i}),$ (5)
$c_{i}$ $=\pi_{\text{c}}(q_{i},\mathcal{H}_{<i}),$
$p_{i}$ $=\pi_{\text{s}}((q_{i},c_{i}),p_{i-1},\mathcal{H}_{<i}).$
By attending to prior reasoning steps, the system forms coherent, memory-informed trajectories across iterations.
Joint Optimization Objective.
We define the multi-agent policy as $\Pi=\{\pi_{\text{t}},\pi_{\text{c}},\pi_{\text{s}}\}$, and optimize it jointly to maximize task performance while ensuring interpretability and alignment with sub-goals:
$\max_{\Pi}\,\mathbb{E}_{(x,y)\sim D}\left[\mathcal{R}(\Pi)-\lambda\sum_{i=1}^{n}\mathcal{L}_{\text{align}}((q_{i},c_{i}),st_{i})\right],$ (6)
where $\mathcal{R}(\Pi)$ denotes the cumulative reward from the Target agent, and $\mathcal{L}_{\text{align}}$ penalizes semantic drift from intended optimization goals.
This tri-agent structure enables interpretable, step-wise refinement of prompts through structured reasoning and localized feedback, offering both flexibility and transparency in discrete prompt optimization.
Proposition 1 (Socratic Policy Improvement Bound). Let $\Pi=\{\pi_{\text{t}},\pi_{\text{c}},\pi_{\text{s}}\}$ denote the joint policy, and suppose the Socratic signal $a_{i}=(q_{i},c_{i})$ induces expected advantage $\bar{A}_{i}>0$ over the prior state $s_{i-1}$. Then, under bounded variance $\sigma^{2}$, the cumulative improvement over $n$ steps satisfies:
$\mathbb{E}[\mathcal{R}(p_{n})]-\mathcal{R}(p_{0})\geq\sum_{i=1}^{n}\left(\bar{A}_{i}-\frac{\sigma^{2}}{2\lambda}\right),$ (7)
where $\lambda$ is the local Lipschitz constant of the reward surface.
This provides a lower bound on improvement, formally linking guidance signal quality to reward trajectory.
Derivation is in Appendix A.1.
@@
Algorithm 1 MARS Optimization Procedure
1:Input: Dataset $\mathcal{D}$, initial prompt $p_{0}$, threshold $\delta$, max iterations II
@@
3:Planner: Generate sub-goal trajectory $\mathbf{ST}=\{st_{1},\dots,st_{n}\}$
4:Initialize $p_{0}^{(1)}\leftarrow p_{0}$, $\mathcal{R}^{(0)}\leftarrow 0$
5:for iteration $t=1$ to II do
6:  for step $i=1$ to $n$ do // Generate question
7:   Teacher generates question $q_{i}\leftarrow\pi_{\text{t}}(st_{i},p_{i-1}^{(t)})$
8:   repeat
9:     Critic evaluates $q_{i}$ & returns feedback $c_{i}\leftarrow\pi_{\text{c}}(q_{i})$
10:     Teacher revises $q_{i}$ if $c_{i}$ is unsatisfactory
11:   until Socratic quality is satisfied
12:   Set $a_{i}\leftarrow(q_{i},c_{i})$
13:   Student updates $p_{i}^{(t)}\leftarrow\pi_{\text{s}}(a_{i},p_{i-1}^{(t)})$
@@
15:  Let $p_{\ell}^{(t)}\leftarrow p_{n}^{(t)}$ // Final prompt
16:  Target evaluates reward
17:  $\mathcal{R}^{(t)}=\sum_{(x,y)\in D_{\text{test}}}f(\pi_{\text{tar}}(x;p_{\ell}^{(t)}),y)$
18:  if $\mathcal{R}^{(t)}-\mathcal{R}^{(t-1)}<\delta$ then
19:   break //Early stopping
20:  end if
21:end for
22:return $p^{*}\leftarrow p_{\ell}^{(t)}$
Evaluation and Iteration
Upon completing the Socratic refinement trajectory, the final prompt $p_{\ell}=p_{n}$—produced through successive dialogue transitions from latent state $s_{0}$ to $s_{\ell}$—is evaluated by the Target agent $\pi_{\text{tar}}$ on the held-out test set $D_{\text{test}}$. The evaluation provides an external signal to measure the effectiveness of the entire optimization trajectory:
$\mathcal{R}^{(t)}=\sum_{(x,y)\in D_{\text{test}}}f\left(\pi_{\text{tar}}(x;p_{\ell}^{(t)}),y\right),$ (8)
where $f(\cdot)$ is a task-specific scoring function (e.g., accuracy, BLEU, F1), and $p_{\ell}^{(t)}$ denotes the final prompt obtained at iteration $t$. This scalar reward serves as the global performance metric, closing the loop between prompt generation and task-level effectiveness.
Adaptive Termination.
To ensure efficient convergence and prevent over-refinement, we adopt an adaptive early stopping criterion based on marginal reward improvement. The gain between two consecutive iterations is defined as:
$\Delta\mathcal{R}^{(t)}=\mathcal{R}^{(t)}-\mathcal{R}^{(t-1)}.$ (9)
The refinement continues only if:
$\Delta\mathcal{R}^{(t)}>\delta,\quad t<I,$ (10)
where $\delta$ is a minimum improvement threshold, and $I$ is the maximum number of allowed iterations.
This iterative control mechanism enforces a form of performance-aware policy halting under the POMDP framework. It ensures that MARS focuses on high-impact updates while avoiding excessive computation on marginally beneficial revisions. As a result, the system adaptively determines the optimal stopping point based on observable task performance.
Proposition 2 (Monotonic Reward Stability). Assume $\mathcal{R}(p)$ is $\lambda$-Lipschitz and each step satisfies $\|p_{i}-p_{i-1}\|\leq\varepsilon$. Then the reward trajectory $\{\mathcal{R}(p_{i})\}$ satisfies:
$|\mathcal{R}(p_{i+1})-\mathcal{R}(p_{i})|\leq\lambda\varepsilon.$
In particular, if $\mathcal{R}(p_{i})<\mathcal{R}(p_{i+1})$ for some $i$, then improvement is bounded and monotonic. This result guarantees bounded gain/loss and motivates early stopping under stable improvement.
Derivation is in Appendix A.2.
Models B.E D.QA $F.$ G.S. R.N. S.U. C.B. C.M. $E.$ W.H. H.A. M.T. Avg.
@@
Experiments
In this section, we present extensive experiments conducted on 12 general task datasets and 5 domain-specific datasets. We begin by introducing the datasets and hyperparameters, followed by the main experimental results. A detailed analysis of the efficiency of the proposed framework is also provided.
More detailed descriptions of the tasks and datasets can be found in Appendix B. The abbreviations used for the tasks in Tables 1 and 2, along with their full names and dataset descriptions, are provided in Appendix B. Baseline methods and additional experimental details are introduced in Appendix C.
Experimental Settings
Datasets.
We select a total of 17 datasets covering both general-purpose and domain-specific tasks. Specifically, we use 6 tasks from the Big-Bench Hard (BBH) suite (suzgun2022challenging) and 6 tasks from MMLU (wang2024mmlu) to represent general reasoning and knowledge-intensive benchmarks. For domain-specific evaluation, we include 3 Chinese subject-area tasks from C-Eval (huang2024c), 1 legal reasoning task from LSAT-AR (zhong2023agieval), and 1 arithmetic reasoning task from GSM8K (zhang2024careful).
Hyperparameters and Evaluation Protocol.
We adopt deepseek-v2.5-1210 (guo2025deepseek) as the primary backbone LLM for all APO tasks. The generation temperature is set to 0.6 to balance creativity and coherence. We configure the maximum number of optimization iterations as $I=10$, with an early stopping threshold of $\delta=0.01$ based on accuracy improvement. To enhance efficiency, each assess-adjust cycle is limited to a single revision per step. Final evaluation is performed using accuracy, computed by comparing the model prediction $y_{\text{pred}}$ with the ground truth label $y$.
Main Results
MARS enhances the average performance across diverse task types.